# Finding local integrals of motion in quantum lattice models in the thermodynamic limit

### v2

@@ I Introduction
Closed macroscopic systems during time-evolution reach a stationary state that is determined by local conservation laws or, equivalently, by local integrals of motion (LIOMs). Generic systems have only a few conserved quantities (like Hamiltonian, particle number, magnetization) and evolve to a thermal Gibbs state [1, 2, 3]. Interacting integrable systems have extensive number of LIOMs and the relevant stationary states have the form that is consistent with the generalized Gibbs ensemble [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. Identification of LIOMs is also the starting point for construction of the hydrodynamics or the generalized hydrodynamics [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31] in case of integrable models. Moreover, anomalous transport (or its anticipated absence) in various quantum systems has recently been linked to the presence of LIOMs. Most prominent examples concern strongly disordered systems [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43], tilted chains with emergent conservation of the dipole moment [44] and systems exhibiting the Hilbert space fragmentation [45, 46, 44].
@@ II.1 Algorithm
Our main assumption is that the dynamics is set either by a local Hamiltonian, $\hat{H}$, or by a repeated sequence of local unitary transformations $\hat{U}$. In the former case, we take $\hat{H}=\sum_{l}\hat{H}_{l}$, where $\hat{H}_{l}$ acts on $M_{H}$ sites. We look for local operators, $\hat{Q}$, which commute with the Hamiltonian, $[\hat{Q},\hat{H}]=0$ or, equivalently, we require vanishing of the norm of a Hermitian operator $\hat{Q}^{\prime}=i[\hat{Q},\hat{H}]$, i.e., $\lVert\hat{Q}^{\prime}\lVert^{2}=\langle\hat{Q}^{\prime}\hat{Q}^{\prime}\rangle=0$. We note that $\hat{Q}^{\prime}$ is still local, $\hat{Q}^{\prime}=\sum_{l}\hat{Q}^{\prime}_{l}$, where $\hat{Q}^{\prime}_{l}$ acts on up to $M+M_{H}-1$ sites. For this reason, $\langle\hat{Q}^{\prime}\hat{Q}^{\prime}\rangle$ can also be calculated exactly in the thermodynamic limit. In case of circuits built out of repeated sequence of local unitary transformations, the same reasoning can be applied for $\hat{Q}^{\prime}=\hat{U}^{\dagger}\hat{Q}\hat{U}-\hat{Q}$. I.e., $\hat{Q}^{\prime}$ is a local operator and vanishing norm of $\hat{Q}^{\prime}$ implies that $\hat{Q}$ is conserved. Both cases can be comprehended within a unified formulation. Namely, we introduce a function $f(\hat{O}^{s})$ which has the following properties: (i) $f$ is linear, (ii) vanishing of $\lVert f(\hat{O}^{s})\lVert$ is equivalent to conservation of $\hat{O}^{s}$; (iii) $f(\hat{O}^{s})$ is a local and Hermitian operator so that one can exactly calculate $\langle f(\hat{O}^{s})f(\hat{O}^{p})\rangle$. If the dynamics is set by a Hamiltonian then one may choose $f(\hat{O}^{s})=i[\hat{O}^{s},\hat{H}]$, whereas $f(\hat{O}^{s})=\hat{U}^{\dagger}\hat{O}^{s}\hat{U}-\hat{O}^{s}$ can be used for unitary circuits. The specific choice of $f$ is irrelevant and can be adapted to the desired setup, as long as it is consistent with conditions (i)-(iii).
@@
Solving the eigenproblem in (1) gives not only the LIOMs but also the Mazur bounds for all operators $\hat{O}^{s}$ used in the construction of the $F$ matrix. Using Eq.(2) one finds that the infinite-time ($t$) correlation functions are bounded by the matrix $V$,
$\lim_{T\to\infty}\frac{1}{T}\int_{0}^{T}{\rm d}t\,\langle\hat{O}^{s}(t)\hat{O}^{s}\rangle\geq\sum_{\alpha=1}^{N_{L}}\frac{\langle\hat{O}^{s}\hat{Q}^{\alpha}\rangle^{2}}{\langle\hat{Q}^{\alpha}\hat{Q}^{\alpha}\rangle}=\sum_{\alpha=1}^{N_{L}}V^{2}_{\alpha s}\,.$ (6)
For systems where the dynamics is set by the Hamiltonian, our approach is consistent with the idea from Ref. [69, 70, 71] based on minimizing the norm of commutators $[\hat{H},\hat{A}]$. For quantum circuits, it is similar (but not equivalent) to the approach from Refs. [72, 73, 74, 75] based on Ruelle-Pollicott resonances obtained from diagonalization of a non-Hermitian matrix $\langle f(\hat{O}^{s})\hat{O}^{p}\rangle$.
@@ II.2 The basis of operators
In the following, we assume that the first, $\hat{\sigma}^{s_{1}}_{l}$, and the last, $\hat{\sigma}^{s_{m}}_{l+m-1}$, spin operators are other than $\hat{\sigma}^{0}$. All operators with maximal support $M$ (i.e., generated for all $m\leq M$) are orthogonal $\langle\hat{O}^{s}_{l}\hat{O}^{s^{\prime}}_{l^{\prime}}\rangle=\delta_{l,l^{\prime}}\delta_{s,s^{\prime}}/L$. For fixed $M$, one has in total $N_{O}=3\times 4^{M-1}$ extensive operators which are orthonormal $\langle\hat{O}^{s}\hat{O}^{s^{\prime}}\rangle=\delta_{s,s^{\prime}}$. We note that all translationally invariant operators supported on up to $M$ sites can be expressed as linear combinations of $\hat{O}^{s}$, thus one may easily calculate the HS product of such operators. Finally, we recall that product of two $\hat{\sigma}^{s}_{l}$-operators is proportional to other such operator, $\hat{\sigma}^{s}_{l}\hat{\sigma}^{s^{\prime}}_{l}=\alpha_{s^{\prime\prime}}\hat{\sigma}^{s^{\prime\prime}}_{l}$, where $\alpha_{s^{\prime\prime}}\in\{1,\pm i\}$. This enables a straightforward calculation of $f(\hat{O}^{s})$, which we implement using the software package from Ref. [76].
@@ III.2 Quasilocal conserved operators and spin stiffness
Figure 3: Three smallest eigenvalues, $\lambda_{\alpha}$, and eigenvectors, $V_{\alpha,1}$, versus maximal support, $M$, for XXZ model within symmetry resolved basis with (a),(b): $\Delta=3/4$ and (c),(d),$\Delta=0.1$.
@@ III.3 Integrable unitary circuit
For every LIOM $\hat{Q}_{n}$ from continuous dynamics, we obtain two LIOMs $\hat{Q}_{n}^{\pm}(\delta)$ in discrete dynamics, related to each other by single-site translation [19]. Both $\hat{Q}_{n}^{\pm}(\delta)$ are supported on $12n+$ 1 sites, and in the limit $\delta\to 0$ reproduce XXX LIOMs. This observation allows us to choose an appropriate operator basis for our algorithm. We start with the same local basis as in the XXZ chain case, but consider extensive operators invariant under two-site translations, $\hat{O}^{{s},-}=\sum_{l}\hat{O}^{{s}}_{2l-1}$ and $\hat{O}^{{s},+}=\sum_{l}\hat{O}^{{s}}_{2l}$. Due to the staggered structure of the propagator $\hat{\mathcal{U}}(\delta)$, we have to relax the assumption that the first operator in the support is different from $\hat{\sigma}^{0}_{l}$. We avoid trivial LIOMs related to magnetization conservation by removing $m=1$ operators from the basis.
@@ III.4 Nearly integrable spin ladders
where $\hat{\sigma}^{s}_{l,y}$ is the spin operator at site $l$ in leg $y=1,2$. This model is integrable for $U=0$ when it represents two decoupled XXZ chains, but also for $\Delta=0$ when it represents the Bose-Hubbard model. We have chosen this Hamiltonian because various relaxation times in the nearly integrable regime ($0<\Delta\ll 1$ and $0<U\ll 1$) may differ by a few orders of magnitude and some of them are extremely long [79]. For this reason, numerical studies of the time evolution are very challenging.
@@
Figure 8: (a) Cross-section of results from Fig. 7 along the red lines shown in Fig. 7. $\Gamma^{\pm}$ shows the relaxation rate of $\hat{Q}^{\pm}=Q^{3}_{y=1}\pm Q^{3}_{2=1}$ estimated from Lanczos propagation method in Ref. [79].
@@ IV Concluding remarks
We have derived and implemented a general numerical approach that provides exact translationally-invariant LIOMs and the corresponding Mazur bounds for infinite quantum systems for which the dynamics is set by either local Hamiltonian or by local unitary transformations. We can obtain the exact LIOMs for the thermodynamic limit because our approach, similarly to Refs. [69, 70, 71], is targeting the short-time dynamics (as encoded, e.g., in the commutators $[\hat{O},\hat{H}]$) under which local operator $\hat{O}$ remains local. Consequently, our results have an important advantage over approaches focused on the infinite-time properties, which are encoded in the diagonal matrix elements of observables calculated in the basis of the Hamiltonian’s eigenstates. In the latter case, the thermodynamic limit is unreachable, and finite-size effects seem unavoidable.

### v3

@@ I Introduction
Closed macroscopic quantum systems undergo a unitary evolution. However, a reduced density matrix of a subsystem is expected to reach a stationary state. The latter is determined by local conservation laws or, equivalently, by local integrals of motion (LIOMs). In particular, the expectation values of local observables are equal to averages over an appropriate ensemble [1]. Generic systems have only a few conserved quantities (like Hamiltonian, particle number, magnetization) and the relevant ensemble has the form of the thermal Gibbs state [2, 3, 4]. Interacting integrable systems have extensive number of LIOMs and the stationary states (of subsystems) have the form that is consistent with the generalized Gibbs ensemble [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. Identification of LIOMs is also the starting point for construction of the hydrodynamics or the generalized hydrodynamics [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32] in case of integrable models. Moreover, anomalous transport (or its anticipated absence) in various quantum systems has recently been linked to the presence of LIOMs. Most prominent examples concern strongly disordered systems [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44], tilted chains with emergent conservation of the dipole moment [45] and systems exhibiting the Hilbert space fragmentation [46, 47, 45].
@@ II.1 Algorithm
Our main assumption is that the dynamics is set either by a local Hamiltonian, $\hat{H}$, or by a repeated sequence of local unitary transformations $\hat{U}$. In the former case, we take $\hat{H}=\sum_{l}\hat{H}_{l}$, where $\hat{H}_{l}$ acts on $M_{H}$ sites. We look for local operators, $\hat{Q}$, which commute with the Hamiltonian, $\hat{Q},\hat{H}]=0$ or, equivalently, we require vanishing of the norm of a Hermitian operator $\hat{Q}^{\prime}=i[\hat{Q},\hat{H}$], i.e., $\lVert\hat{Q}^{\prime}\lVert^{2}=\langle\hat{Q}^{\prime}\hat{Q}^{\prime}\rangle=0$. We note that $\hat{Q}^{\prime}$ is still local, $\hat{Q}^{\prime}=\sum_{l}\hat{Q}^{\prime}_{l}$, where $\hat{Q}^{\prime}_{l}$ acts on up to $M+M_{H}-1$ sites. For this reason, $\langle\hat{Q}^{\prime}\hat{Q}^{\prime}\rangle$ can also be calculated exactly in the thermodynamic limit. In case of circuits built out of repeated sequence of local unitary transformations, the same reasoning can be applied for $\hat{Q}^{\prime}=\hat{U}^{\dagger}\hat{Q}\hat{U}-\hat{Q}$. I.e., $\hat{Q}^{\prime}$ is a local operator and vanishing norm of $\hat{Q}^{\prime}$ implies that $\hat{Q}$ is conserved. Both cases can be comprehended within a unified formulation. Namely, we introduce a function $f(\hat{O}^{s})$ which has the following properties: (i) $f$ is linear, (ii) vanishing of $\lVert f(\hat{O}^{s})\lVert$ is equivalent to conservation of $\hat{O}^{s}$; (iii) $f(\hat{O}^{s})$ is a local and Hermitian operator so that one can exactly calculate $\langle f(\hat{O}^{s})f(\hat{O}^{p})\rangle$. If the dynamics is set by a Hamiltonian then one may choose $f(\hat{O}^{s})=i[\hat{O}^{s},\hat{H}$], whereas $f(\hat{O}^{s})=\hat{U}^{\dagger}\hat{O}^{s}\hat{U}-\hat{O}^{s}$ can be used for unitary circuits. The specific choice of $f$ is irrelevant and can be adapted to the desired setup, as long as it is consistent with conditions (i)-(iii).
@@
In what follows we demonstrate that the eigenproblem in (1) gives not only the LIOMs but also the Mazur bounds for all operators $\hat{O}^{s}$ used in the construction of the $F$ matrix. To this end we start from a standard form of the Mazur bound [69]
$\lim_{T\to\infty}\frac{1}{T}\int_{0}^{T}{\rm d}t\,\langle\hat{O}^{s}(t)\hat{O}^{s}\rangle\geq\sum_{\alpha=1}^{N_{L}}\frac{\langle\hat{O}^{s}\hat{Q}^{\alpha}\rangle^{2}}{\langle\hat{Q}^{\alpha}\hat{Q}^{\alpha}\rangle},$ (6)
and note that all $\hat{Q}^{\alpha}$ are orthonormal, $\langle\hat{Q}^{\alpha^{\prime}}\hat{Q}^{\alpha}\rangle=\delta_{\alpha,\alpha^{\prime}}$, so are the operators $\hat{O}^{s}$, $\langle\hat{O}^{s}\hat{O}^{s^{\prime}}\rangle=\delta_{s,s^{\prime}}$. Then, using Eq. (2) one may evaluate the expression on the right-hand side of Eq. (6)
$\sum_{\alpha=1}^{N_{L}}\frac{\langle\hat{O}^{s}\hat{Q}^{\alpha}\rangle^{2}}{\langle\hat{Q}^{\alpha}\hat{Q}^{\alpha}\rangle}=\sum_{\alpha=1}^{N_{L}}\sum_{s^{\prime}}V^{2}_{\alpha s^{\prime}}\langle\hat{O}^{s}\hat{O}^{s^{\prime}}\rangle^{2}=\sum_{\alpha=1}^{N_{L}}V^{2}_{\alpha s}.\$(7)
For systems where the dynamics is set by the Hamiltonian, our approach is consistent with the idea from Ref. [70, 71, 72] based on minimizing the norm of commutators $\hat{H},\hat{A}$]. For quantum circuits, it is similar (but not equivalent) to the approach from Refs. [73, 74, 75, 76] based on Ruelle-Pollicott resonances obtained from diagonalization of a non-Hermitian matrix $\langle f(\hat{O}^{s})\hat{O}^{p}\rangle$.
@@ II.2 The basis of operators
In the following, we assume that the first, $\hat{\sigma}^{s_{1}}_{l}$, and the last, $\hat{\sigma}^{s_{m}}_{l+m-1}$, spin operators are other than $\hat{\sigma}^{0}$. All operators with maximal support $M$ (i.e., generated for all $m\leq M$) are orthogonal $\langle\hat{O}^{s}_{l}\hat{O}^{s^{\prime}}_{l^{\prime}}\rangle=\delta_{l,l^{\prime}}\delta_{s,s^{\prime}}/L$. For fixed $M$, one can generate an exhaustive list by considering $\hat{O}^{s}_{l}$ with the first operator chosen from the restricted set $\{\hat{\sigma}^{x}_{l},\hat{\sigma}^{y}_{l},\hat{\sigma}^{z}_{l}\}$, and subsequent operators from the full set, including the identity. This guarantees no duplicates are present in the basis set after all the $N_{O}=3\times 4^{M-1}$ extensive operators are built. They are orthonormal $\langle\hat{O}^{s}\hat{O}^{s^{\prime}}\rangle=\delta_{s,s^{\prime}}$ by construction. We note that all translationally invariant operators supported on up to $M$ sites can be expressed as linear combinations of $\hat{O}^{s}$, thus one may easily calculate the HS product of such operators. Finally, we recall that product of two $\hat{\sigma}^{s}_{l}$-operators is proportional to other such operator, $\hat{\sigma}^{s}_{l}\hat{\sigma}^{s^{\prime}}_{l}=\alpha_{s^{\prime\prime}}\hat{\sigma}^{s^{\prime\prime}}_{l}$, where $\alpha_{s^{\prime\prime}}\in\{1,\pm i\}$. This enables a straightforward calculation of $f(\hat{O}^{s})$, which we implement using the software package from Ref. [77].
@@ III.2 Quasilocal conserved operators and spin stiffness
[…]
@@ III.3 Integrable unitary circuit
For every LIOM $\hat{Q}_{n}$ from continuous dynamics, we obtain two LIOMs $\hat{Q}_{n}^{\pm}(\delta)$ in discrete dynamics, related to each other by single-site translation [20]. Both $\hat{Q}_{n}^{\pm}(\delta)$ are supported on $2n+1$ sites, and in the limit $\delta\to 0$ reproduce XXX LIOMs. This observation allows us to choose an appropriate operator basis for our algorithm. We start with the same local basis as in the XXZ chain case, but consider extensive operators invariant under two-site translations, $\hat{O}^{{s},-}=\sum_{l}\hat{O}^{{s}}_{2l-1}$ and $\hat{O}^{{s},+}=\sum_{l}\hat{O}^{{s}}_{2l}$. Due to the staggered structure of the propagator $\hat{\mathcal{U}}(\delta)$, we have to relax the assumption that the first operator in the support is different from $\hat{\sigma}^{0}_{l}$. We avoid trivial LIOMs related to magnetization conservation by removing $m=1$ operators from the basis.
@@ III.4 Nearly integrable spin ladders
where $\hat{\sigma}^{s}_{l,y}$ is the spin operator at site $l$ in leg $y=1,2$. This model is integrable for $U=0$ when it represents two decoupled XXZ chains, but also for $\Delta=0$ when it represents the Fermi-Hubbard model. We have chosen this Hamiltonian because various relaxation times in the nearly integrable regime ($0<\Delta\ll 1$ and $0<U\ll 1$) may differ by a few orders of magnitude and some of them are extremely long [80]. For this reason, numerical studies of the time evolution are very challenging.
@@
Figure 8: Cross-section of results from Fig. 7 along the red lines shown in Fig. 7. $\Gamma^{\pm}$ shows the relaxation rate of $\hat{Q}^{\pm}=Q^{3}_{y=1}\pm Q^{3}_{2=1}$ estimated from Lanczos propagation method in Ref. [80].
@@ IV Concluding remarks
We have derived and implemented a general numerical approach that provides exact translationally-invariant LIOMs and the corresponding Mazur bounds for infinite quantum systems for which the dynamics is set by either local Hamiltonian or by local unitary transformations. We can obtain the exact LIOMs for the thermodynamic limit because our approach, similarly to Refs. [70, 71, 72], is targeting the short-time dynamics (as encoded, e.g., in the commutators $\hat{O},\hat{H}$]) under which local operator $\hat{O}$ remains local. Consequently, our results have an important advantage over approaches focused on the infinite-time properties, which are encoded in the diagonal matrix elements of observables calculated in the basis of the Hamiltonian’s eigenstates. In the latter case, the thermodynamic limit is unreachable, and finite-size effects seem unavoidable.